40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1411     Entry time: Fri May 6 12:38:49 2011
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: re-modematched and realigned, locked 

Last night I calculated new modematching solutions for both beams using the profiles measured yesterday morning. The REFL beams now return to the FIs at a manageably small size (very close to the size of the beams coming through from the input side, as they should). See the picture below. Note that both beams are incident on the card: input from one side and REFL from the other. It was impossible to notice a difference in size by eye---this is good.


I went about realigning the cavity mirrors over again, since it seems like they might have been bumped around with all the commotion. I then locked the cavity in the CCW direction and fine-tuned to get about 50% transmission. We should be able to get about 90% with reasonable losses and our R values, but we've never seemed to be able to get much higher than half. The modematching solutions I calculated should give us the standard >99% coupling via the naïve model, but I still maintain that we can't get a truly flat phasefront with an astigmatic input beam, and can therefore not get the coupling we calculate with 'EtaCalcSimple'.


I was nearly done with aligning the CW beam when something happened and the CCW refused to stay locked. Things were also difficult to work with because the output optics were no longer aligned so that the two output beams were overlapping, so I couldn't simultaneously monitor both while aligning stuff. The best way to do this is to reconfigure all that output stuff the way we used to have it and then go about optimizing both beam alignments. Working on this today. 

ELOG V3.1.3-